OpenStudy (anonymous):

3n^2-n/n^2-1 divided by n^2/n+1

OpenStudy (anonymous):

Refer to the attachment.

jhonyy9 (jhonyy9):

step by step : 3n^2 -n = n(3n-1) n^2 -1 = (n-1)(n+1) hope you know again that a --- b a d ---- = --- * ---- c b c --- d hope so much that theses will help you